OpenStudy (anonymous):

Two new hybrid vehicles accelerate at different rates. The Ultra Range's acceleration can be modeled by the function d(t) = 1.5t^2, while the Edisons can be modeled by the function d(t) = 5.4t^2, where d is the distance, in meters, and t is the time in seconds. The ultra Range starts the race at 0s. At what time should the Edison start so that both cars are at the same point 5s after the race starts? Express your answers to the nearest tenth of a second.

OpenStudy (anonymous):

d(t) = 1.5t^2 U d(t) = 5.4t^2 E 1.5(5)^2= 5.4t^2 sft 2.6352313

OpenStudy (anonymous):

What does sft 2.63 mean?

OpenStudy (anonymous):

solve for t :) t is 2.63

OpenStudy (anonymous):

the Edison car should start at 2.63 to catch up to the ultra

OpenStudy (anonymous):

The answer in the book says 2.4 seconds

OpenStudy (anonymous):

idk then that is odd

OpenStudy (anonymous):

no you are correct timo86m all you have to do next is subtract 5-2.635231..=2.4. "At what time should the edison start so that both cars are at the same point at 5" so you must subtract find the diffrence and that will give you your answer.